# An improved perturbative QCD study of the decays $B_{c}^{+}\to\eta_{c}L^{+}$

### v2

@@ III.1 Factorizable-emission dominated $B_{c}^{+}\to\eta_{c}L^{+}$
Next, for the $B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}$ channel, its BR in the iPQCD formalism is read as,
@@
The future precise measurements would provide useful information to help study the QCD dynamics among the states pion, $\rho$ and $a_{1}(1260)$ involving the same quark components.
To provide referenced values for experimental measurements of this mode, the $B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}(\to\pi^{+}\pi^{-}\pi^{+})$ and $B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}(\to K^{+}K^{-}\pi^{+})$ BRs could be derived in the iPQCD formalism under the narrow-width approximation via resonance state $a_{1}(1260)$ with the strong decays ${\cal B}(a_{1}(1260)^{+}\to\pi^{+}\pi^{-}\pi^{+})=0.50\pm 0.05$ Aubert:2007kpb and ${\cal B}(a_{1}(1260)^{+}\to K^{+}K^{-}\pi^{+})=0.11\pm 0.02$ Liu:2023kxr as,
${\rm BR}(B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}(\to\pi^{+}\pi^{-}\pi^{+}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+})\cdot{\cal B}(a_{1}^{+}\to\pi^{+}\pi^{-}\pi^{+})$ (54)
$=$ $(3.46^{+1.04}_{-0.86})\times 10^{-3}\;,$
and
${\rm BR}(B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}(\to K^{+}K^{-}\pi^{+}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+})\cdot{\cal B}(a_{1}^{+}\to K^{+}K^{-}\pi^{+})$ (55)
$=$ $(7.60^{+2.57}_{-2.21})\times 10^{-4}\;,$
In principle, these large BRs around ${\cal O}(10^{-3})$ could be accessed easily in the related experiments. Moreover, interestingly, Ref. Wang:2012vna ever provided the $B_{c}^{+}\to\eta_{c}\pi^{+}\pi^{-}\pi^{+}$ BR as $1.854\times 10^{-3}$ with $B_{c}\to\eta_{c}$ form factor calculated by QCD sum rules, associated with the relevant ratios between the BRs of $B_{c}^{+}\to\eta_{c}\pi^{+}\pi^{-}\pi^{+}$ and $B_{c}^{+}\to(J/\psi,\eta_{c})\pi^{+}$ as $1.52$ and $1.36$, respectively. For comparison, these values can also be given in the iPQCD formalism,
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+}\pi^{-}\pi^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $=$ $2.96^{+1.20}_{-0.99}\;,\qquad\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+}\pi^{-}\pi^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}=1.70^{+0.68}_{-0.55}\;,$ (56)
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K^{+}K^{-}\pi^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $=$ $0.65^{+0.28}_{-0.23}\;,\qquad\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K^{+}K^{-}\pi^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}=0.37^{+0.16}_{-0.13}\;.$ (57)
These numerical values could be confronted with the future measurements.
@@
where the 1st (2nd) entry in Eqs. (60) and (63) corresponds to $\theta_{K}=33^{\circ}\ (58^{\circ})$. The similar patterns to the BRs also appear in the following observables for related modes trivially.
@@
In the above numerical results, the dominant errors arise from the uncertainties of shape parameter $\beta_{B_{c}}$ and Gegenbauer moments in the distribution amplitudes of $B_{c}$, and $K_{1A}$ and $K_{1B}$ states, respectively. The $B_{c}^{+}\to\eta_{c}K_{1}(1270,1400)^{+}$ BRs indicate that ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+})$ is consistent well with ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+})$ at $\theta_{K}\sim 33^{\circ}$ within uncertainties, while ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+})$ is significantly larger than ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+})$ at $\theta_{K}\sim 58^{\circ}$ with a factor around 6. It means that an destructive interference between $B_{c}^{+}\to\eta_{c}K_{1A}^{+}$ and $B_{c}^{+}\to\eta_{c}K_{1B}^{+}$ occurs significantly in $B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+}$ at $\theta_{K}\sim 58^{\circ}$. In other words, future measurements testing these $B_{c}^{+}\to\eta_{c}K_{1}(1270,1400)^{+}$ BRs with clearly different results would lead to a better determination of $\theta_{K}$. A verified $\theta_{K}$ value is key to further guide the theoretical predictions with good precision.
For the future probe of $B_{c}^{+}\to\eta_{c}K_{1}(1270,1400)^{+}$ at LHC experiments, by inputting the data ${\cal B}(K_{1}(1270)^{+}\to K^{+}\rho^{0}(\to\pi^{\pm}\pi^{\mp}))=0.190\pm 0.065$ and ${\cal B}(K_{1}(1400)^{+}\to K^{*0}(\to K^{\pm}\pi^{\mp})\pi^{+})=0.313\pm 0.020$ ParticleDataGroup:2024cfk , the BRs of multibody channels $B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+}(\to K^{+}\pi^{\pm}\pi^{\mp})$ and $B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+}(\to K^{\pm}\pi^{\mp}\pi^{+})$ via resonances $K_{1}(1270,1400)$ could be deduced in the iPQCD formalism under the narrow-width approximation,
${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+}(\to K^{+}\pi^{\pm}\pi^{\mp}))$ $=$ $\left\{\begin{array}[]{ll}(4.88^{+2.50}_{-2.31})\times 10^{-5}&\vskip 3.41418pt\\ (7.60^{+3.51}_{-3.25})\times 10^{-5}&\end{array}\right$., (71)
${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+}(\to K^{\pm}\pi^{\mp}\pi^{+}))$ $=$ $\left\{\begin{array}[]{ll}(6.57^{+2.48}_{-2.02})\times 10^{-5}&\vskip 3.41418pt\\ (2.10^{+1.23}_{-0.89})\times 10^{-5}&\end{array}\right$., (74)
@@
It is worth emphasizing that, the determination of $\theta_{K}$ with definite value is highly important, because, if $\theta_{K}$ could be determined unambiguously, it could further help constrain the mixing between $f_{1}(1285)(h_{1}(1170))$ and $f_{1}(1420)(h_{1}(1450))$ promisingly Cheng:2011pb . The ratios between the $B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+}$ and $B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+}$ BRs are then derived to provide necessary reference for constraining the magnitude of $\theta_{K}$,
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+})}$ $=$ $\left\{\begin{array}[]{ll}0.82^{+0.09}_{-0.08}&\vskip 3.41418pt\\ 0.17^{+0.04}_{-0.04}&\\ \end{array}\right$., (77)
@@
The above BRs and the associated ratios predicted in the iPQCD formalism would be helpful to explore the QCD dynamics in the considered axial-vectors, especially in $K_{1}(1270,1400)$.
III.2 Factorizable-emission suppressed $B_{c}^{+}\to\eta_{c}L^{+}$
Now, we turn to analyze the factorizable-emission suppressed $B_{c}^{+}\to\eta_{c}L^{+}$ decays, including $B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+}$, $B_{c}^{+}\to\eta_{c}S^{+}$, and $B_{c}^{+}\to\eta_{c}T^{+}$, respectively. In explicit words, different to the factorizable-emission dominated $B_{c}^{+}\to\eta_{c}(P,V)^{+}$ decays, the factorizable-emission contributions in the former two kinds of channels are generally suppressed due to the tiny zeroth Gegenbauer moment $a_{0,b_{1}}^{\parallel}$ in the longitudinal leading-twist $b_{1}(1235)^{+}$ distribution amplitude Yang:2005gk and the nearly-zero vector decay constant $f_{S}$ for light scalars, respectively, while the latter one with even vanished factorizable-emission amplitudes at leading order is just because of the fact that a tensor meson cannot be produced via vector currents.
For the decay $B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+}$, its iPQCD BR is read as,
accompanied by the following two ratios,
Notice that, the BRs of $B_{c}^{+}\to(J/\psi,\eta_{c})b_{1}(1235)^{+}$ in the iPQCD formalism indicate an interesting relation, i.e., ${\rm BR}(B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+})\simeq{\rm BR}(B_{c}^{+}\to J/\psi b_{1}(1235)^{+})\sim{\cal O}(10^{-3})$ within uncertainties, even the latter BR containing three kinds of polarization contributions. These values could be tested in the near-future experiments to help understand the QCD dynamics in the related decays, as well as in $b_{1}(1235)$.
According to the categorization of light scalar mesons in two scenarios, the CP-averaged $B_{c}^{+}\to\eta_{c}S^{+}$ BRs in the iPQCD formalism are given as,
@@ III.2 Factorizable-emission suppressed $B_{c}^{+}\to\eta_{c}L^{+}$
where the 1st (2nd) entry in Eqs. (84) and (87) corresponds to $S1(S2)$ and the dominant errors come mainly from the $B_{c}$-meson shape parameter $\beta_{B_{c}}$, and from the scalar decay constant $\bar{f}_{S}$ and the Gegenbauer moments $B_{i}(i=1,3)$ of scalar mesons, respectively. The future precise constraints from experimental measurements and Lattice QCD calculations on these hadronic parameters are urgently demanded for theoretical predictions with good precision.
@@
It seems that the errors in these results induced by the hadronic parameters are indeed cancelled to a great extent. However, it is surprisingly noted that, drastically different to the values presented in Eqs. (40) and (50), the above two ratios are significantly large near ${\cal O}(10^{2})$ within uncertainties, even already with a known factor of $|V_{us}/V_{ud}|^{2}\sim 0.05$. Moreover, they are also remarkably larger than those in the $B_{c}^{+}\to J/\psi S^{+}$ decays correspondingly within the iPQCD framework, for detail, see equation (61) in Liu:2023kxr .
Table 4: Decay amplitudes (in units of $10^{-3}$GeV-3) of $B_{c}^{+}\to\eta_{c}S^{+}$ from different twists in the iPQCD formalism. The upper (lower) entry corresponds to the scalars $a_{0}(1450)^{+}$ and ${K_{0}^{*}}(1430)^{+}$ in scenario 1 (2) at every line. For simplicity, only the central values are quoted for clarifications.
@@
$\phi_{\eta_{c}}^{v}(x)$ $\phi_{\eta_{c}}^{s}(x)$ $\phi_{\eta_{c}}^{v}(x)$ $\phi_{\eta_{c}}^{s}(x)$
$B_{c}^{+}\to\eta_{c}a_{0}(980)^{+}$ $0.57-{\it i}1.65$ $0.25-{\it i}0.23$ $62+{\it i}2.80 0.0$$03-{\it i}0.04$
$B_{c}^{+}\to\eta_{c}\kappa^{+}$ $08-{\it i}14.79 5.5$$54+{\it i}24.88$ $1.76-{\it i}2.28$ $0.25-{\it i}0.35$
$B_{c}^{+}\to\eta_{c}a_{0}(1450)^{+}$ $\begin{array}[]{cc}-0.08-{\it i}0.21\\ -0.67-{\it i}1.91\end{array}$ $\begin{array}[]{cc}0.10+{\it i}0.34\\ 0.76+{\it i}3.32\end{array}$ $\begin{array}[]{cc}-0.15+{\it i}0.06\\ 0.20-{\it i}0.03\end{array}$ $\begin{array}[]{cc}0.01+{\it i}0.02\\ -0.03-{\it i}0.02\end{array}$
$B_{c}^{+}\to\eta_{c}{K_{0}^{*}}(1430)^{+}$ $\begin{array}[]{cc}-0.97-{\it i}2.59\\ -5.26-{\it i}15.07\end{array}$ $\begin{array}[]{cc}1.18+{\it i}4.19\\ 5.97+{\it i}26.08\end{array}$ $\begin{array}[]{cc}-1.17+{\it i}0.55\\ 1.63-{\it i}0.39\end{array}$ $\begin{array}[]{cc}0.13+{\it i}0.13\\ -0.22-{\it i}0.14\end{array}$

### v3

@@ III.1 $B_{c}^{+}\to\eta_{c}(P,V)^{+}$
Table 2: Various predictions for ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$, ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$ and ${\rm BR}(B_{c}^{+}\to\eta_{c}K^{+})/{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ in this work and in the literature, respectively.
Observables This work Chang:1992pt Anisimov:1998uk Colangelo:1999zn AbdElHady:1999xh Verma:2001hb Ebert:2003cn Ivanov:2006ni Hernandez:2006gt Choi:2009ym  [^1] The results are calculated on the basis of Coulomb plus linear confining and harmonic oscillator (in the parentheses) potentials. Naimuddin:2012dy Qiao:2012hp  [^2] These results are calculated with the non-relativistic QCD approach at leading order in the strong coupling $\alpha_{s}$. And the corresponding next-to-leading-order BR in $\alpha_{s}^{2}$ is $(5.19^{+0.70}_{-1.07})\times 10^{-3}$. Rui:2014tpa Issadykov:2018myx Nayak:2022qaq Wu:2024gcq Deng:2025znr  [^3] These results are calculated with the QCD factorization approach in $\alpha_{s}$. And the corresponding next-to-next-to-leading-order BR in $\alpha_{s}^{2}$ is $(0.81^{+0.88}_{-0.56})\times 10^{-3}$.
$10^{3}\cdot{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ $2.03^{+0.53}_{-0.41}$ $1.80$ $1.30$ $0.26$ $1.40$ $9.30$ $0.85$ $1.90$ $0.94$ $0.91(1.16)$ $0.34$ $2.95$ $2.98^{+1.24}_{-1.05}$ $1.89^{+0.37}_{-0.37}$ $0.397$ $1.40^{+0.40}_{-0.40}$ $0.79^{+0.86}_{-0.55}$
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $1.74^{+0.66}_{-0.50}$ $1.00$ $1.78$ $0.19$ $1.27$ $2.07$ $1.39$ $1.12$ $1.24$ $\dots$ $1.00$ $1.33$ $1.28^{+0.69}_{-0.56}$ $1.87^{+0.52}_{-0.52}$ $1.02$ $0.56^{+0.21}_{-0.21}$ $0.95^{+1.04}_{-0.68}$
$10^{2}\cdot R^{\eta_{c}}_{K/\pi}$ $49^{+0.54}_{-0.49} 7.7$87.78 $10.00$ $8.00$ $7.86$ $5.05$ $8.24$ $7.89$ $7.98$ $8.13(8.10)$ $8.82$ $7.12$ $8.05^{+0.68}_{-0.99}$ $7.94^{+0.02}_{-0.05}$ $7.81$ $7.86^{+0.14}_{-0.08}$ $7.51^{+0.02}_{-0.00}$
Table 3: Same as Table 2 but for $B_{c}^{+}\to\eta_{c}(\rho,K^{*})^{+}$.
Observables This work Chang:1992pt Anisimov:1998uk Colangelo:1999zn AbdElHady:1999xh Verma:2001hb Ebert:2003cn Ivanov:2006ni Hernandez:2006gt Choi:2009ym Naimuddin:2012dy Qiao:2012hp Rui:2014tpa Issadykov:2018myx Nayak:2022qaq Wu:2024gcq Deng:2025znr
$10^{3}\cdot{\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})$ $5.37^{+1.46}_{-1.12}$ $4.90$ $3.00$ $0.67$ $3.30$ $370.0$ $2.10$ $4.50$ $2.40$ $2.57(3.24)$ $1.06$ $7.89$ $9.83^{+3.33}_{-2.59}$ $5.18^{+1.04}_{-1.04}$ $1.24$ $3.80^{+0.10}_{-0.10}$ $2.15^{+2.23}_{-1.46}$
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $4.59^{+1.77}_{-1.34}$ $2.72$ $4.11$ $0.52$ $3.0$ $82.2$ $3.44$ $2.65$ $3.16$ $\dots$ $3.12$ $3.55$ $4.22^{+2.05}_{-1.57}$ $5.13^{+1.45}_{-1.45}$ $3.18$ $1.52^{+0.37}_{-0.37}$ $2.60^{+2.71}_{-1.80}$
$10^{2}\cdot R^{\eta_{c}}_{K^{*}/\rho}$ $5.68^{+0.11}_{-0.11}$ $5.10$ $7.00$ $5.97$ $5.45$ $3.70$ $5.24$ $5.56$ $5.42$ $5.06(5.25)$ $5.66$ $5.20$ $5.80^{+0.52}_{-0.45}$ $5.60^{+0.03}_{-0.04}$ $5.24$ $5.53^{+1.14}_{-1.48}$ $5.12^{+0.15}_{-0.06}$
Finally, as aforementioned in the Introduction, the decays $B_{c}^{+}\to\eta_{c}(P,V)^{+}$ have been studied extensively with various $B_{c}\to\eta_{c}$ form factors in different approaches Chang:1992pt ; Anisimov:1998uk ; Colangelo:1999zn ; AbdElHady:1999xh ; Verma:2001hb ; Ebert:2003cn ; Ivanov:2006ni ; Hernandez:2006gt ; Choi:2009ym ; Naimuddin:2012dy ; Qiao:2012hp ; Rui:2014tpa ; Issadykov:2018myx ; Nayak:2022qaq ; Wu:2024gcq ; Deng:2025znr . However, the obtained values of BRs and relative ratios have a wide spread, as explicitly presented in Tables 2 and 3. Based on the numerical results in the literature and this work, some remarks are in order: (a) For the absolute BRs of $B_{c}^{+}\to\eta_{c}(\pi,\rho)^{+}$, as seen in the first line of Tables 2 and 3, it is clear that, within $1\sigma$ errors, our iPQCD predictions are consistent with those in Refs. Chang:1992pt ; Ivanov:2006ni ; Rui:2014tpa ; Issadykov:2018myx ; Wu:2024gcq ; Deng:2025znr for ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ and with those in Refs. Chang:1992pt ; Ivanov:2006ni ; Issadykov:2018myx ; Deng:2025znr for ${\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})$, respectively. We can also easily find that the results of BRs for both $B_{c}^{+}\to\eta_{c}\pi^{+}$ and $B_{c}^{+}\to\eta_{c}\rho^{+}$ decays given in Refs. Colangelo:1999zn ; Verma:2001hb ; Ebert:2003cn ; Hernandez:2006gt ; Naimuddin:2012dy ; Nayak:2022qaq are inconsistent with ours, even within $2\sigma$ errors. Needless to say, there are significant discrepancies in the understanding of the transition $B_{c}\to\eta_{c}$ across different frameworks. (b) For the relative ratios ${\rm BR}(B_{c}^{+}\to\eta_{c}(\pi,\rho)^{+})/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$, as collected in the second line of Tables 2 and 3, it is found that, within $1\sigma$ errors, our iPQCD predictions are consistent with those in Refs. Anisimov:1998uk ; AbdElHady:1999xh ; Verma:2001hb ; Ebert:2003cn ; Hernandez:2006gt ; Qiao:2012hp ; Rui:2014tpa ; Issadykov:2018myx ; Deng:2025znr for ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$ and with those in Refs. Anisimov:1998uk ; Ebert:2003cn ; Qiao:2012hp ; Rui:2014tpa ; Issadykov:2018myx ; Deng:2025znr for ${\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$, respectively. (c) As shown in the third line of Tables II and III, within $1\sigma$ errors, our predictions for $R^{\eta_{c}}_{K/\pi}$ are consistent with most literature results except those in Refs. Anisimov:1998uk ; Verma:2001hb ; Ebert:2003cn ; Naimuddin:2012dy , while for $R^{\eta_{c}}_{K^{*}/\rho}$ they agree only with those in Refs. AbdElHady:1999xh ; Ivanov:2006ni ; Naimuddin:2012dy ; Rui:2014tpa ; Issadykov:2018myx ; Wu:2024gcq .
It is worth noting that all our numerical results, such as individual BRs and various relative ratios, for the decays $B_{c}^{+}\to\eta_{c}(P,V)^{+}$ in the iPQCD formalism agree well with those calculated in covariant confined quark model Issadykov:2018myx within $1\sigma$ uncertainties. To evaluate the reliability of the available theoretical calculations across different models, future precise measurements are highly essential.
III.2 $B_{c}^{+}\to\eta_{c}(A,S,T)^{+}$
In this subsection, we present the predictions of the decays $B_{c}^{+}\to\eta_{c}(A,S,T)^{+}$ for the first time within the QCD-based factorization framework, which could provide references for understanding the structure of these considered light hadrons in the future.
First of all, for the $B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}$ channel, its BR in the iPQCD formalism is read as
@@ III.2 $B_{c}^{+}\to\eta_{c}(A,S,T)^{+}$
In the above results, the uncertainties are mainly dominated by the variations of shape parameter $\beta_{B_{c}}$ in the $B_{c}$-meson distribution amplitude. Future precise measurements will provide valuable information to aid in the study of QCD dynamics among the states pion, $\rho$, and $a_{1}(1260)$, which involve the same quark components.
For the decay $B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+}$, in principle, different from the $B_{c}^{+}\to\eta_{c}(P,V)^{+}$ modes, its factorizable-emission contributions are generally suppressed due to the tiny zeroth Gegenbauer moment $a_{0,b_{1}}^{\parallel}$ in the longitudinal leading-twist $b_{1}(1235)^{+}$ distribution amplitude Yang:2005gk , while its iPQCD BR is read as
accompanied by the two ratios
in which theoretical uncertainties of these numerical results are dominantly induced by the variations of shape parameter $\beta_{B_{c}}$ and Gegenbauer moment $a_{b_{1}}$ in the leading-twist distribution amplitudes of $B_{c}$ and $b_{1}(1235)$, respectively. Notice that, the BRs of $B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+}$ and $B_{c}^{+}\to J/\psi b_{1}(1235)^{+}$ in the iPQCD formalism indicate an interesting relation, namely, ${\rm BR}(B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+})\simeq{\rm BR}(B_{c}^{+}\to J/\psi b_{1}(1235)^{+})\sim{\cal O}(10^{-3})$ within uncertainties, even though the latter BR includes three kinds of polarization contributions Liu:2023kxr . These values could be tested in the near-future experiments to help understand the QCD dynamics in the related decays, as well as in $b_{1}(1235)$.
@@
where the first (second) entry in Eqs. (58) and (61) corresponds to $\theta_{K}=33^{\circ}\ (58^{\circ})$. The similar patterns to the BRs also appear in the following observables for related modes trivially:
@@
In the above numerical results, the dominant errors arise from the uncertainties of shape parameter $\beta_{B_{c}}$ and Gegenbauer moments in the distribution amplitudes of $B_{c}$, and $K_{1A}$ and $K_{1B}$ states, respectively. The $B_{c}^{+}\to\eta_{c}K_{1}(1270,1400)^{+}$ BRs indicate that ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+})$ is very consistent with ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+})$ at $\theta_{K}\sim 33^{\circ}$ within uncertainties, while ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+})$ is significantly larger than ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+})$ at $\theta_{K}\sim 58^{\circ}$ with a factor around 6. It means that a destructive interference between $B_{c}^{+}\to\eta_{c}K_{1A}^{+}$ and $B_{c}^{+}\to\eta_{c}K_{1B}^{+}$ occurs significantly in $B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+}$ at $\theta_{K}\sim 58^{\circ}$. In other words, future measurements testing these $B_{c}^{+}\to\eta_{c}K_{1}(1270,1400)^{+}$ BRs with clearly different results would lead to a better determination of $\theta_{K}$. A verified $\theta_{K}$ value is key to further guide the theoretical predictions with good precision.
@@
It is worth emphasizing that the determination of $\theta_{K}$ with definite value is highly important because, if $\theta_{K}$ could be determined unambiguously, it could further help constrain the mixing between $f_{1}(1285)(h_{1}(1170))$ and $f_{1}(1420)(h_{1}(1450))$ promisingly Cheng:2011pb . The ratios between the $B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+}$ and $B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+}$ BRs are then derived to provide necessary reference for constraining the magnitude of $\theta_{K}$,
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+})}$ $=$ $\left\{\begin{array}[]{ll}0.82^{+0.09}_{-0.08}&\vskip 3.41418pt\\ 0.17^{+0.04}_{-0.04}&\\ \end{array}\right.$. (69)
@@
The above BRs and the associated ratios predicted in the iPQCD formalism would be helpful to explore the QCD dynamics in the considered axial vectors, especially in $K_{1}(1270,1400)$.
Next, we turn to the $B_{c}^{+}\to\eta_{c}S^{+}$ decays. As stressed in the Introduction, relative to the above $B_{c}^{+}\to\eta_{c}(P,V)^{+}$ decays, the nearly zero vector decay constant $f_{S}$ for light scalars will lead to the significantly suppressed factorizable-emission contributions in the $B_{c}^{+}\to\eta_{c}S^{+}$ channels. According to the categorization of light scalar mesons in two scenarios, the CP-averaged $B_{c}^{+}\to\eta_{c}S^{+}$ BRs in the iPQCD formalism are given as
@@
where the first (second) entry in Eqs. (74) and (77) corresponds to $S1(S2)$ and the dominant errors come mainly from the $B_{c}$-meson shape parameter $\beta_{B_{c}}$ and from the scalar decay constant $\bar{f}_{S}$ and the Gegenbauer moments $B_{i}(i=1,3)$ of scalar mesons, respectively. The future precise constraints from experimental measurements and Lattice QCD calculations on these hadronic parameters are urgently demanded for theoretical predictions with good precision. Nevertheless, since the LHC experiments can measure the $B_{c}$ decays with a branching ratio at $10^{-6}$ level Descotes-Genon:2009eui , then the BRs $(7.20^{+3.69}_{-2.62})\times 10^{-6}$ for $B_{c}^{+}\to\eta_{c}\kappa^{+}$ and $(6.68^{+3.81}_{-2.50})\times 10^{-7}[(1.29^{+0.91}_{-0.58})\times 10^{-5}]$ for $B_{c}^{+}\to\eta_{c}{K_{0}^{*}}(1430)^{+}$ are expected to be probed in the newly upgraded LHCb experiment. Furthermore, the experimental confirmations on the evident order difference of magnitudes in the predicted BRs of $B_{c}^{+}\to\eta_{c}{K_{0}^{*}}(1430)^{+}$ could potentially provide important clues to ascertain the preferred scenario for understanding the scalar $K_{0}^{*}(1430)$.
@@
It seems that the errors in these results induced by the hadronic parameters are indeed canceled to a great extent. However, it is surprisingly noted that, drastically different from the values presented in Eqs. (40) and (50), the above two ratios are significantly large near ${\cal O}(10^{2})$ within uncertainties, even though already with a known factor of $|V_{us}/V_{ud}|^{2}\sim 0.05$. Moreover, they are also remarkably larger than those in the $B_{c}^{+}\to J/\psi S^{+}$ decays correspondingly within the iPQCD framework, for detail, see Eq. (61) in Ref. Liu:2023kxr .
Table 4: Decay amplitudes (in units of $10^{-3}$GeV-3) of $B_{c}^{+}\to\eta_{c}\;(J/\psi)\;S^{+}$ in the iPQCD formalism. The upper (lower) entry corresponds to the scalars $a_{0}(1450)^{+}$ and ${K_{0}^{*}}(1430)^{+}$ in scenario 1 (2) at every line. For simplicity, only the central values are quoted for clarifications.
@@
$B_{c}^{+}\to\eta_{c}\;(J/\psi)\;a_{0}(980)^{+}$ $0.32-{\it i}1.88\;(-0.11-{\it i}0.73)$ $0.59+{\it i}2.76\;(-48.13+{\it i}52.70)$
$B_{c}^{+}\to\eta_{c}\;(J/\psi)\;\kappa^{+}$ $0.46+{\it i}10.09\;(-1.17-{\it i}7.60)$ $1.51-{\it i}2.63\;(-7.87+{\it i}12.70)$
$B_{c}^{+}\to\eta_{c}\;(J/\psi)\;a_{0}(1450)^{+}$ $\begin{array}[]{cc}0.02+{\it i}0.13\;(0.09+{\it i}0.63)\\ 0.09+{\it i}1.41\;(-0.15-{\it i}0.97)\end{array}$ $\begin{array}[]{cc}-0.14+{\it i}0.08\;(-68.32+{\it i}33.43)\\ 0.17-{\it i}0.05\;(-44.16+{\it i}7.08)\end{array}$
$B_{c}^{+}\to\eta_{c}\;(J/\psi)\;{K_{0}^{*}}(1430)^{+}$ $\begin{array}[]{cc}0.21+{\it i}1.60\;(1.07+{\it i}7.06)\\ 0.71+{\it i}11.01\;(-1.60-{\it i}10.60)\end{array}$ $\begin{array}[]{cc}-1.04+{\it i}0.68\;(-11.39+{\it i}6.47)\\ 1.41-{\it i}0.53\;(-10.30+{\it i}2.90)\end{array}$